OpenStudy (anonymous):

The sum of the measures of two exterior angles of a triangle is 260. What is the measure of the third exterior angle? Can you show me how to do this?

OpenStudy (anonymous):

The exterior angles of a triangle will always add up to 360.

OpenStudy (anonymous):

ah i see thanks

OpenStudy (anonymous):

No worries :)

OpenStudy (anonymous):

So the third would be 100?

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

 A+B+C=180 A+a+B+b+C+c=180+180+180=540 a+b+c=540-(A+B+C)=360 a+b=260, given c=100